p1=10kN
need members BA,BE, and DE.
2 m + 2 m KON ON S SA TE VN 1 1.5 m SA
Answer #1

JokN tend=2 / x[Q = 53.13) lism *am * grant with poina C Efyzo Desind=0 DC70) {fazo Dccolo + CD o arazo) Doint D lOKN Ecoc {f

- point B E Am ľ se (a=57199 ar la En Becond + BD + Be caldo 0+ (-10) + BE COM57:17) - 2 BE=16.667 KN Efnod -BA - BE sind + B

Hence, membe Hence, member member force. I Force (KM) -17.32 KN DE 16,667KN DE TO (-ve) a compresian (the) Tension

Feel free to ask for further clarification.

please up-vote if you liked this solution. As it will encourage me to solve more and more problems and helping students.

Thank you,